OpenStudy (anonymous):

What is the measure of the angle formed by the minute hand and the hour hand of a clock when the clock shows 4:35? A. 72.5 b.90 c. 120 d. 72

OpenStudy (anonymous):



OpenStudy (anonymous):

120?

OpenStudy (anonymous):

i don't know grab a clock and measure it

OpenStudy (anonymous):

really? if your gonna be like that maybe you shouldn't write to people that need help

OpenStudy (anonymous):

alright you're right, i'm sorry that definitely looks like 120, put that in for your answer

OpenStudy (anonymous):

Alright. So use 12: 00 as the origin point. Because you're taking both measurements from there, you can subtract them, no problem. Think of an hour. You're 35 minutes through one, so what proportion of the clock have you gone through? It should be a little bit over half, right? Because you're just over halfway through an hour. So you're a little over halfway through 360 degrees. So the angle of the minute hand should be a little over 180, right? So what's that number? Do the same thing with the hour hand. You're comparing 4 hours to 12 hours, so what's the proportion, and how many degrees out of 360 is that?

OpenStudy (anonymous):

Minute Hand:\[\frac{1}{3}2\pi +\frac{35}{60}\frac{\pi }{12}=\frac{103 \pi }{144}=128.75{}^{\circ} \]Hour Hand:\[ \frac{35}{60}360=210{}^{\circ} \]Angles measured clockwise.

OpenStudy (anonymous):

My bad. The descriptions should be reversed.

OpenStudy (anonymous):

Actually, yeah, it does depend on whether the hour hand adjusts like a real clock. In that case, you're not looking at 4 out of 12, but 4 and then the fraction of minutes out of an hour that it is past the four o'clock mark.